# AI-Augmented Thyroid Scintigraphy for Robust Classification

### v1

@@ 6 Conclusion
This study examined the impact of various augmentation strategies, including SD-, FM-, and CA-based methods, on thyroid classification using ResNet18. Our results show that FM-based augmentation, particularly when combined with the original dataset (O+FM) or CA (O+FM+CA), consistently outperformed SD-based approaches. FM enables smoother, more controlled image transformations, preserving key structural and intensity details crucial for classification.
{credits}
6.0.1 Acknowledgements
This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant RGPIN-2019-06467.
[…]
The authors have no competing interests to declare.

### v2

@@ 6 Conclusion
In this study, we explored the impact of different augmentation strategies, including SD-, FM-, and CA-based, on thyroid classification using ResNet18. Our findings demonstrated that FM-based augmentation, particularly when combined with the original dataset (O+FM), consistently led to superior performance compared to SD-based approaches. The addition of CA (O+FM+CA) further improved classification accuracy and ensured more balanced performance across thyroid diagnostic groups (DG, NG, NL, and TI).
Statistical significance testing using the Wilcoxon method reinforced these results, highlighting the effectiveness of FM in improving model generalization. FM ensures a smoother and more controlled transformation of image distributions, preserving essential structural details and intensity variations critical for classification. In contrast, SD-based models, especially those relying on prompt-only (SD2), image-only (SD3), or masked inputs (SD4), may introduce artifacts or inconsistencies that can mislead the classifier. This enhanced realism in FM-generated images leads to better feature representation learning, improving classification performance.
Overall, this work presents a methodology enabling more objective and consistent thyroid scintigraphy analysis, and therefore diagnosis, with the potential to assist Nuclear Medicine physicians and ultimately help improve patient outcomes.